Problem 2 D small manufacturing company is manufact has prring company is planing production requirements requ ected demand bplanning production requirements for next year. The Sales by quarter as shown below: Winter Average quarterly demand 3500 units 3800 units 4200 units 4500 units Suppose the company had 50 workers and 1.600 units in inventory at the beginning of the winter The company wishes to finish the year with an inventory of 1,000 units on hand. The average pay per worker is $3,000 per month, including benefits. Production per worker is 50 s per month. It costs $600 to hire a new worker and $1,500 to lay off a worker. Inventory carrying cost averages $40 per quarter per unit. Using the table below, determine the total cost of using Level strategy Resources Winter Spring Summer Fall Total Beg. Inventory Required # of workers Production Demand Ending Inventory Costs Regular labor cost Hiring/firing cost Inv. Carrying cost Other costs, if any Total
Answer #1

Net demand = Sum of demand of four quarters + Desired ending inventory - Beginning inventory

= 4200+4500+3500+3800+1000-1600

= 15400

Average demand per season = 15400/4 = 3850

Number of level workforce required = 3850/50 = 77

Regular labor cost per unit = Average pay per worker per month / Production per worker per month = $ 3000/50 = $ 60

Aggregate plan based on Level strategy is following:

1 2 Aggregate plan based on Level Strategy Resources Beg. Invento 4 Number of Workers50 Winter Sprin Summer Fall Total 4,400 308 15,400 16,000 3800 1,600 1,250 600 950 5 Production 6 Demand 7 Ending invento 3,850 4,200 1,250 3,850 4,500 600 3,850 3,500 950 3,850 3,800 1,000 1,600 9 Costs 10 Regular labor cost$60231,000 231,000231,000 231,000 924,000 16,200 152,000 1Hiring / Firing cost 12 Inv. carrying cost 13 Other costs, if an 14 Total 15 16,200 50,000 0 0 0 S40 24,000 38,000 40,000 297,200255,000 269,000271,000 1,092,200 17 20 21

EXCEL FORMULAS:

1 Aggregate plan based on Level Strategy 2 Resources Winter Summer B7 -D7 4 Number of Workers 50 5 Production 6 Demand 7 Ending inven -C4 -D4 50 4500 -D4 C4 50 4200 -C3+C5-C6 E4 50 3500 F4 50 3800 -F3+F5-F6 1600 D3+D5-D6 E3+E5-E6 9 Costs 10 Regular labor cost 3000/50SB10*C5 11 Hiring/ Firing cost 12 Inv. carrying cost 40 13 Other costs, if an 14 Total 15 -$B10 E5 MAX(0,E4-D4) 600+MAXI0,D4-E4) 1500 -MAX(0,F4-E4 $B12 E7 $B10 F5 -MAX(0,C4-B4) 600+MAX(0,B4-C4) 1500 -SB12 c7 -SB10 D5 MAX(0,D4-C4) 600+MAX(0,C4-D4) 1500 -SB12 D7 SUM(C10:C13) SUM(D10:D13) SUM(E10:E13) SUM(F10:F1 17 182 3 -D7 4 -D4 Total SUM(C3:F3 ESUM(C4:F4) SUM(C5:F5 SUM(C6:F6) SUM(C7:F7) Summer Fall -F4 50 3800 -F3+F5-F6 5 E4 50 6 3500 10 SB10*E5 11 MAX(0,E4-D4)*600+MAX(0,D4-E4)1500 MAX(0,F4-E4)*600+MAX(0,E4-F4) 1500 SUM(C11:F11) 12 SB12 E7 -$B10 *F5 SUM(C10:F10) -SB12*F7 SUM(C12:F12 SUM(C13:F13) SUM(C14:F14) 14 ESUM(E10:E13) 15 16 17 SUM(F10:F13 20 21